OpenStudy (anonymous):

Solve on interval 0,2pi

OpenStudy (anonymous):

\[(cosx+1)(2\cos^2 x-3cosx-2)\]

OpenStudy (anonymous):

ok hold on

hartnn (hartnn):

there's no equation to solve :O what you've written is an expression, is the Q by any chance, \((cosx+1)(2\cos^2 x-3cosx-2)=0\) ?

OpenStudy (anonymous):

\[x=2\pi,x=\frac{ \pi }{ 2 },x=\frac{ \pi }{ 3 }\] \[x=\frac{ \pi }{ 6 },x=\frac{ 7\pi }{ 6 }\] \[x=\pi,x=\frac{ 2\pi }{ 3 },x=\frac{ 4\pi }{ 3 }\] \[x=2\pi,x=\frac{ \pi }{ 2 },x=\frac{ 5\pi }{ 4 }\]

OpenStudy (anonymous):

no sorry thats all i have

OpenStudy (anonymous):

@hartnn help?

hartnn (hartnn):

its this only : \((cosx+1)(2\cos^2 x-3cosx-2)=0\) so, we have \((cosx+1)=0 \quad and \quad (2\cos^2 x-3cosx-2)=0\) the first one is easy, any ideas what to do for \( (2\cos^2 x-3cosx-2)=0\) ?

OpenStudy (anonymous):

simplifies to \[(2cosx+1)(cosx-2)\] ?

OpenStudy (anonymous):

graph it lol.

hartnn (hartnn):

yes, \((2cosx+1)(cosx-2)=0\) is correct. now just solve these, \((2cosx+1)=0 \quad \\ (cosx-2)=0\)

OpenStudy (anonymous):

how would i sole them?

hartnn (hartnn):

2 cos x +1 =0 isolate cos x, can u ?

hartnn (hartnn):

like for cos x -2 = 0 cos x = 2

OpenStudy (anonymous):

\[cosx=-\frac{ 1 }{ 2 }\]?

hartnn (hartnn):

thats correct, now for which values of x, is cos x = -1/2 ?

OpenStudy (anonymous):

what do you mean? sorry ive been up all night!!!

hartnn (hartnn):

i mean there are 2 angles for which cos x equals -1/2 like, 2 values for which cos x =+1/2 will be pi/3 and 5pi/3

hartnn (hartnn):

also, cos x = 2 is not possible because cos x can take values from +1 to -1 only.

OpenStudy (anonymous):

how would i find the radical?

hartnn (hartnn):

from the unit circle, you can find that , cos x = -1/2 for x =2pi/3 and 4pi/3

OpenStudy (anonymous):

okay

hartnn (hartnn):

and cos x +1 =0 would give you cos x = -1 you know for which angle is cos x = -1 ??

OpenStudy (anonymous):

do you have an example of a unit circle? i found one online but its confusing

hartnn (hartnn):

you just look where, x-cordinate (cos) equals, -1/2 if you search, you will find it at 2 point, x = 2pi/3 and x= 4pi/3 now search where x-co-ordinate is -1 ?

OpenStudy (anonymous):

\[\pi,\frac{ 3\pi }{ 2 }?\]

hartnn (hartnn):

not at 3pi/2 at 3pi/2, x-cordinate = 0 so, only x= pi

OpenStudy (anonymous):

oops sorry

hartnn (hartnn):

that means cos pi=-1

hartnn (hartnn):

now we have only 3 roots , x=pi, 2pi/3,4pi/3

OpenStudy (anonymous):

so it would be C?

hartnn (hartnn):

yes :) any doubts ?

OpenStudy (anonymous):

Yay! you're so smart! :)

hartnn (hartnn):

you too are smart enough :)

OpenStudy (anonymous):

lol smart enough! :) gee thanks!

hartnn (hartnn):

welcome ^_^